OpenStudy (anonymous):

Need help! Tell whether the given number is a solution of each equation. 1. 5x + 10 = -35; -9 and 5(-9) + 10 = -35...

OpenStudy (anonymous):

5(-9)+10=-35 -45+10=-35 -35=-35

OpenStudy (anonymous):

Thanks. but. I still dont understand.. My book has both problems list and then it ask the question.. Tell whether the given number is a solution of each equation..